OpenStudy (anonymous):

Solve for x (0≤x≤2π) 4sin2xcos2x−sin2x−(√3)cos2x=0

OpenStudy (dumbcow):

i found 8 solutions: \[x = \frac{\pi}{9}, \frac{\pi}{6}, \frac{4\pi}{9}, \frac{7\pi}{9}, \frac{10\pi}{9}, \frac{7\pi}{6}, \frac{13\pi}{9}, \frac{16\pi}{9}\] not sure how to solve it using trig identities though, sorry

OpenStudy (anonymous):

That hould be right because there are 8 solutions. But I really need the steps. I've simplified it to \[4\sin2x - \tan2x - \sqrt{3} = 0\] Does that help?